OpenStudy (anonymous):

4x-3/5=2/5 solve for x~!

OpenStudy (anonymous):

4x-(3)/(5)=(2)/(5) Since -(3)/(5) does not contain the variable to solve for, move it to the right-hand side of the equation by adding (3)/(5) to both sides. 4x=(3)/(5)+(2)/(5) Simplify the right-hand side of the equation. 4x=1 Divide each term in the equation by 4. (4x)/(4)=(1)/(4) Simplify the left-hand side of the equation by canceling the common factors. x=1/4